# Inverse Problem Sampling in Latent Space Using Sequential Monte Carlo

### v1

@@ 5.3 Analysis
Recall that in our proposal update we suggested correcting the prior mean using two terms, one that involves ${\mathbf{y}}_{0}$ and another that involves ${\mathbf{y}}_{t}$. In practice during the sampling process, we first use the former one and then switch to the latter at some fixed time step $s$. This is a hyperparameter of our approach. Table 3 compares the FID and PSNR on ImageNet inpainting task as a function of the diffusion step $s$ in which the switch is made. From the table, when the switch is done at earlier stages of the sampling process, the FID improves, but at the same time, the PSNR degrades. We chose to use $s=333$ since it balances well between the two metrics while giving more emphasis to the perceptual quality.
Table 3: Tradeoff between the proposal update correction terms for $T=1000$ steps.
FID ($\downarrow$) PSNR ($\uparrow$)
$s=0$ 65.76 19.61
$s=T/6$ 61.09 19.25
$s=T/3$ 51.81 18.87
$s=T/2$ 48.77 18.59
6 Limitations
Although our approach has strong empirical results, one limitation of our approach is related to computational demand. The sampling time and the memory demand increase with the number of particles. In addition, compared to TDS, in the resampling step, we need to use the decoder one more time to compute $p({\mathbf{y}}_{t}|{\mathbf{z}}_{t})$ (only forward pass), which can also affect the sampling time. This effect can be mitigated by taking fewer particles or by parallelizing LD-SMC between GPUs. Furthermore, for more challenging tasks like box inpainting, competing algorithms tend to exhibit noticeable artifacts, which limits their applicability.
@@ 7 Conclusion
In this study, we presented LD-SMC, a novel method for solving inverse problems in the latent space of diffusion models using SMC. Specifically, we leveraged the forward process of the diffusion process to augment the model with auxiliary observations, one per each timestep, and used these observations to guide the sampling process as part of the backward diffusion process. This framework can be seen as applying one step of blocked Gibbs sampling. To perform SMC sampling, we suggested a novel weighing scheme and a novel proposal distribution. Both are based on information from the auxiliary labels and the true label ${\mathbf{y}}_{0}$. Empirically, we validated LD-SMC against strong baseline methods on common benchmarks. The results suggest that LD-SMC can improve the performance over baseline methods, especially in cases where extrapolation is needed (e.g., in inpainting).

### v2

@@ 5.3 Analysis
In Figure 5 we evaluate LD-SMC in terms of FID and PSNR on inpainting tasks when varying $s$, the hyperparameter that controls the influence of ${\mathbf{y}}_{t}$ in the proposal distribution, and $N$, the number of particles. From Figure 5(a) there is a clear trade-off in FID and PSNR when $s$ changes. Higher $s$ values result in a better FID at the expense of PSNR, while lower $s$ values show the opposite trend. For inpainting, $s=333$ strikes a good balance between the two metrics while giving more emphasis to the perceptual quality. In addition, note that taking $s=0$ is similar to using the DPS as a proposal distribution which results in a significant reduction in perceptual quality. Figure 5(b) shows an improvement in FID while suffering from a small reduction in PSNR when increasing the number of particles. In our comparisons, we chose $N=5$ to balance between performance and computational cost incurred by adding particles (see Section E for further details). Empirically, we found that sometimes only increasing the number of particles resulted in favorable metric values, as is also clear from the tables. We attribute this to the relatively small number of particles with which we experimented (up to $10$). However, in general, one can expect that as the number of particles increases, the samples will become more accurate. Finally, in Appendix F we present an ablation study on the effect of using several Gibbs steps. We show that improvements in FID and PSNR can be obtained by taking more than one step, yet there is not always a clear trend of improvements over the steps.
@@ 6 Conclusion
In this study, we presented LD-SMC, a novel method for solving inverse problems in the latent space of diffusion models using SMC. LD-SMC augment the model with auxiliary observations, one per each timestep, and use these observations to guide the sampling process as part of the backward diffusion process. This framework can be seen as applying one step of blocked Gibbs sampling. To perform SMC sampling, we suggested a novel weighing scheme and proposal distributions. Both are based on information from the auxiliary labels and the true label ${\mathbf{y}}_{0}$. Empirically, we validated LD-SMC against strong baseline methods on common benchmarks. The results suggest that LD-SMC can improve the performance over baseline methods, especially in cases where extrapolation is needed (e.g., in inpainting tasks). one limitation of our approach is related to computational demand. The sampling time and the memory demand increase with the number of particles and Gibbs iterations. This effect can be partially mitigated by taking fewer particles and GPU parallelization.